OpenStudy (anonymous):

The length of the rectangle is 3 times the width. the perimeter is 96 cm. finthe the width and length.

OpenStudy (anonymous):

perimeter = 2 ( a + b) also, length is 3x the width so could write: perimeter = 2 ( 3 a + b) 96 cm = 2 ( 3 a + b)